OpenStudy (anonymous):

Solve the following equation: (7x – 9)(3x + 1) = 0

OpenStudy (anonymous):

uh well using the zero-product property, you either have 7x-9 = 0 or 3x+1 = 0 7x - 9 = 0 7x = 9 x = 9/7 3x+ 1 = 0 3x = -1 x = -1/3